# Sequential Monte-Carlo testing by betting

### v4

@@ 7 Experiments on simulated and real data
We begin with three subsections on simulations, and end with two subsections on real data. In the first three subsections, we simulated $m=2000$ independent treatment vs. control trials (see Example 1). For each trial, we generated $n=1000$ observations, where the probability that an observation was treated is $0.5$. Responses of control observations were generated from $N(0,1)$ and responses of treated observations from $N(\mu,1)$. We test the null hypothesis of independence between the treatment and response. For $Y_{t}$, $t\geq 0$, we chose the difference in mean between the treated and untreated observations and $B=1000$ permutations were drawn randomly with replacement.
7.1 Comparison to the permutation p-value
In Figure 1 we compare the (log-transformed) p-values obtained by the classical permutation p-value $\mathfrak{p}^{\mathrm{perm}}_{B}$ with the sequential p-values obtained from the binomial mixture strategy, the binomial strategy and the aggressive strategy. The p-values of the sequential strategies are defined by $1/\max_{b=1,\ldots,B}W_{b}$. Due to Ville’s inequality, these are valid p-values. The binomial mixture strategy was applied with density $u_{c}$, $c=0.9\alpha$, and the parameter of the binomial strategy was chosen as in Algorithm 2.
@@ 7.1 Comparison to the permutation p-value
The results show that the permutation p-value performs best in all cases, except for $\mu=1$, where the aggressive strategy and the permutation p-value are the same. Nevertheless, the number of p-values below the level $\log(\alpha)$, which is illustrated by the grey horizontal line, is nearly the same for the binomial mixture strategy, binomial strategy and the permutation p-value. Therefore, if we test at level $\alpha$, the power should be quite similar, while the number of permutations is lower using the sequential strategies. Further, note that $B$ is fixed for the permutation p-value in advance, while we could decide to carry on testing using the sequential strategies in cases where the current wealth looks promising.
7.2 Comparison to the Besag-Clifford method
In Figure 2 and 3, we explore the power and number of permutations needed to obtain the decision for $\alpha=0.05$ and $\alpha=0.01$. The power is displayed in the upper left plots, the average number of permutations in the upper right plots and the average number of permutations when the testing process was stopped for rejection and futility in the lower left and right plots, respectively. The sequential strategies are stopped for futility if the wealth drops below $\alpha$ and for rejection if the wealth exceeds $1/\alpha$. We replaced the classical permutation p-value $\mathfrak{p}^{\mathrm{perm}}_{B}$ with the Besag-Clifford sequential strategy [3] with parameter $h=\alpha B$. In this case, the power of $\mathfrak{p}^{\mathrm{BC}}_{\gamma(h,B)}$ is the same as the power of $\mathfrak{p}^{\mathrm{perm}}_{B-1}$, even though the Besag-Clifford approach may stop before the maximum number of permutations $B$ are sampled [28].
Figure 2: Power and average number of permutations until the decision was obtained for $\alpha=0.05$ and different strengths of the alternative $\mu$ — following the experimental protocol in Subsection 7.2. Upper left plot: Empirical power; Upper right plot: Average number of permutations $\bar{\tau}$; Lower left plot: Average number of permutations until it was stopped for rejection $\bar{\tau}_{1}$; Lower right plot: Average number of permutations until it was stopped for futility $\bar{\tau}_{0}$. Relationship between the plots is given by $\bar{\tau}=(\bar{\tau_{1}}m_{1}+\bar{\tau}_{0}m_{0}+B(m-m_{1}-m_{0}))/m$, where $m_{1}$ and $m_{0}$ are the number of simulation runs where it was stopped for rejection and futility, respectively. The power of binomial strategy and binomial mixture strategy is similar as with Besag-Clifford and number of permutations is reduced substantially, particularly when the alternative is strong.
@@ 7.3 Improving power by stochastic rounding
Figure 4 shows that stochastic rounding leads to a significant improvement in power, particularly for the binomial mixture strategy in case of $\alpha=0.01$. This also implies that the binomial mixture strategy was often stopped while having a promising wealth. Therefore, drawing $1000$ permutations might be too less when the $\alpha$ is low and the decision is difficult ($\mathfrak{p}^{\mathrm{lim}}$ close to $\alpha$). Hence, the binomial mixture strategy is likely to achieve further rejections if we would continue sampling. In Figure 5 we compare the power of these randomized strategies with the Besag-Clifford method. There is hardly any power difference visible. In particular, the Besag-Clifford method and randomized binomial mixture strategy seem to overlap completely. Note that our methods obtained this power while reducing the number of permutations considerably (Figure 2 and Figure 3) and offering the option to stop at any point in time.
@@ 7.4 Real data: testing Fisher’s sharp null
We consider a treatment vs. control trial with binary outcomes that was already used in [7, 23]. There were 18 successes among the 32 treated observations and 5 successes among the 21 control observations. We are interested in testing Fisher’s sharp null hypothesis of no individual treatment effect $H_{0}^{\mathrm{IT}}$ at level $\alpha=0.05$. Under the Stable Unit Treatment Value Assumption (SUTVA; [24]), which states that there is only one version of the treatment and no interference between the individuals, this null hypothesis can be tested using the permutation test described in Example 1. Fisher’s exact test offers a p-value of $0.026$ and thus the null hypothesis could be rejected by this test. We applied our binomial strategy (Algorithm 2) and binomial mixture strategy with density $u_{c}$, $c=0.95\alpha$, $1000$ times with a maximum number of $B=5000$ permutations. In this case, we omitted the stop for futility to not affect the power. The null hypothesis was rejected by both strategies in all $1000$ runs. The binomial strategy needed a mean number of $85$ permutations and a median number of $53$ permutations, while the binomial mixture strategy required a mean number of $147$ permutations and median number of $97$ permutations to obtain the decision. We illustrated the wealth processes of different runs for the first $500$ steps in Figure 6. Since the binomial mixture strategy converges to $1/c$, we expect it to bet more safely when it approaches $1/\alpha$. For this reason, the binomial strategy allows to reject the hypothesis earlier in most cases. However, when there are a lot of losses, e.g. as in the beginning of the upper right plot, the binomial mixture strategy leads to a larger wealth and thus seeming more promising when the evidence is not very strong. Also note that when the number of wins is much larger than the number of losses, as in the lower plots, the binomial strategy begins to lose wealth when further wins occur — after the hypothesis is already rejected. Due to (7), this is the case if $\mathrm{rank}(Y_{0};Y_{0}^{t-1})/(t+1)<p$, where $p=1/\lceil\sqrt{2\pi e^{1/6}}/\alpha\rceil$.
@@ 7.5 Real data: testing conditional independence under model-X
Here, we consider testing the independence between two variables $W_{0}$ and $X$ conditional on some covariates $Z$, as described in Example 2. We follow the analysis by Grünwald et al. [11] and Berrett et al. [2] and apply our methods on the Capital bikeshare dataset, which is available at https://ride.capitalbikeshare.com/system-data. In this example, $W_{0}$ denotes the logarithm of the ride duration, $X$ the binary variable membership (casual users vs. long-term membership) and $Z$ is the three dimensional vector of starting point, destination and starting time. For the generation of the test statistics $Y_{0},\ldots,Y_{B}$, we used the code of Grünwald et al. [11], which is available in the supplementary material of their paper. The distribution $W_{0}|Z$ is modeled as Gaussian $N(\mu_{Z},\sigma_{Z}^{2})$, where $\mu_{Z}$ and $\sigma_{Z}^{2}$ are estimated by a kernel regression of $W_{0}$ with the time of day as covariate, separately for each combination of starting point and destination. The test statistics are then given by $Y_{i}=|\mathrm{cor}(X,W_{i}-\mathbb{E}[W_{0}|Z])|$.
We applied our binomial strategy and binomial mixture strategy $1000$ times with the same parameters as in Section 7.4 (in particular, the type I error level was set to 0.05). In the original code of Grünwald et al. [11], the training dataset consists of $158,741$ observations and the test data of $7,173$ observations. In this case, all $Y_{i}$, $i\geq 1$, were smaller than $Y_{0}$ such that the binomial strategy needed $44$ permutations and the binomial mixture strategy $61$ permutations to reject the null hypothesis. We also applied our methods when only the first $2000$ observations of the test data were included, thus increasing the probability of $Y_{i}$ being greater or equal than $Y_{0}$. This led to a mean permutation p-value of $0.008$ and the null hypothesis was still rejected by both sequential strategies in all $1000$ runs. The binomial strategy needed a mean number of $49$ permutations and a median number of $44$ permutations, while the binomial mixture strategy required a mean number of $83$ permutations and median number of $61$ permutations to obtain the decision.
@@ 8 Discussion
We have introduced a new test martingale for exchangeability based on a simple testing by betting approach. In contrast to previous works we focused on the alternative that all test statistics are exchangeable except the first. This alternative is motivated by permutation tests, where the observed data is fixed and additional data is generated that is supposed to have the same distribution as the original data under the null hypothesis. Our general betting approach encompasses all rank-based permutation tests and particularly provides anytime-valid generalizations of the classical permutation p-value and the Besag-Clifford method. We derived the log-optimal betting strategy under the considered alternative and showed that the corresponding wealth does not grow exponentially, which distinguishes our setting from many other betting problems. Furthermore, we showed that for any given prior on the limiting permutation p-value $\mathfrak{p}^{\mathrm{lim}}$, the log-optimal strategy can be written as a mixture of simple binomial strategies. Based on this, we proposed a strategy with zero resampling risk after a finite number of permutations for all $p^{\mathrm{lim}}\in[0,c)$, where $c<\alpha$ can be chosen arbitrarily close to $\alpha$. Furthermore, we performed a simulation study and a real data analysis which illustrates that our anytime-valid p-values have a similar power as the classical permutation p-value while the number of permutations until the decision is obtained can be substantially reduced.
@@
Binomial strategy     Reduces permutations substantially     Similar power as classical permutation      p-value in all simulation scenarios     Always rejects earlier than classical      permutation p-value, if      $\mathfrak{p}^{\mathrm{perm}}<1/\lceil\sqrt{2\pi e^{1/6}}/\alpha\rceil\approx\alpha/e$      (Proposition 4.4)     Parameter-free (see Algorithm 2)     Cannot ensure arbitrarily      small resampling risk
@@ Acknowledgments
AR thanks Leila Wehbe for posing the question of whether permutation tests can be stopped early under null and alternative, and for stimulating practical discussions. LF and AR thank Peter Grünwald for a helpful comment. LF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 281474342/GRK2224/2. AR was funded by NSF grant DMS-2310718.
Appendix A More details on (orthogonally) related work
@@
For the particular problem of testing our null hypothesis that the observations are exchangeable, there exists no nontrivial test martingale for the canonical data filtration (generated by the observations) [20]. There are two different approaches to circumvent this issue. The first idea is to replace the data filtration by a coarser one. This means the information to be released at each step is restricted. This was used by Vovk and colleagues in a conformal prediction approach [32, 30], and this is also the approach taken in the current paper (modulo some differences around randomization of ranks). The second approach is to construct an e-process which is not a test martingale, based on universal inference [34, 20]. The idea is to partition the composite null hypothesis, define a test martingale for each subset and take the infimum over these test martingales to yield an e-process. In terms of testing by betting, it could be interpreted as playing multiple games, where the gambler’s wealth is the minimum wealth across all these games. In this case, the wealth process is adapted to the canonical data filtration, has expectation of at most one under the null at any stopping time in the original filtration of the data.
@@
Very recently, other sequential tests of exchangeability were proposed, based on the former approach. Saha and Ramdas [25] introduced a test martingale for Markovian alternatives that is based on “pairwise betting”, which means that two observations are observed and bet on simultaneously. Independently, Koning [16] introduced a sequential test for general group invariance (including exchangeability). In particular, he derives likelihood ratios for group invariance and defines a test martingale as running product of these. Our test martingale is also based on coarsening the data filtration by not revealing the $Y_{t}$ itself, but the rank among the previously known observations, and is thus somewhat related to this line of work. The main difference of our approach to the previously mentioned approaches is the alternative we test against. Vovk [30] focuses on change-point alternatives, Ramdas et al. [20] and Saha and Ramdas [25] on Markovian alternatives, while Koning [16] considers general group invariance versus normal distributions with a location shift. In contrast, our paper focuses on the usual setting of permutation tests (where permutations rather than data are sampled sequentially), where under the alternative, only the original test statistic is special while all the permuted statistics are exchangeable by design.
Appendix B Calibrating $\mathfrak{p}^{\mathrm{perm}}_{B}$ into an e-value

### v5

@@ 7 Experiments on simulated and real data
We begin with three subsections on simulations, and end with two subsections on real data. In the first three subsections, we simulated $m=2000$ independent treatment vs. control trials (see Example 1). For each trial, we generated $n=1000$ observations, where the probability that an observation was treated is $0.5$. Responses of control observations were generated from $N(0,1)$ and responses of treated observations from $N(\mu,1)$. We test the null hypothesis of independence between the treatment and response. For $Y_{t}$, $t\geq 0$, we chose the difference in mean between the treated and untreated observations and $T=1000$ permutations were drawn randomly with replacement.
7.1 Behavior of the anytime-valid p-values
In Figure 1 we compare the (log-transformed) p-values in ascending order obtained by the classical permutation p-value $Q^{\mathrm{perm}}_{T}$ with the sequential p-values obtained from the binomial mixture strategy, the binomial strategy and the aggressive strategy. The p-values of the sequential strategies are defined by $1/\max_{b=1,\ldots,T}W_{b}$. Due to Ville’s inequality, these are valid p-values. The binomial mixture strategy was applied with density $u_{c}$, $c=0.9\alpha$, and the parameter of the binomial strategy was chosen as in Algorithm 2.
@@ 7.1 Behavior of the anytime-valid p-values
The results show that the permutation p-value performs best in all cases, except for $\mu=1$, where the aggressive strategy and the permutation p-value are the same. Nevertheless, the number of log-p-values below the level $\log(\alpha)$, which is illustrated by the grey horizontal line, is nearly the same for the binomial mixture strategy, binomial strategy and the permutation p-value. Therefore, if we test at level $\alpha$, the power should be quite similar, while the number of permutations is lower using the sequential strategies (we will examine the power in detail in the subsequent sections). Further, note that $T$ is fixed for the permutation p-value in advance, while we could decide to carry on testing using the sequential strategies in cases where the current wealth looks promising.
7.2 Power and sample size comparison with the Besag-Clifford method
In Figure 2 and 3, we explore the power and number of permutations needed to obtain the decision for $\alpha=0.05$ and $\alpha=0.01$. The power is displayed in the upper left plots, the average number of permutations in the upper right plots and the average number of permutations when the testing process was stopped for rejection and futility in the lower left and right plots, respectively. The sequential strategies are stopped for futility if the wealth drops below $\alpha$ and for rejection if the wealth exceeds $1/\alpha$. We replaced the classical permutation p-value $Q^{\mathrm{perm}}_{T}$ with the Besag-Clifford sequential strategy [3] with parameter $h=\alpha T$. In this case, the power of $Q^{\mathrm{BC}}_{\gamma(h,T)}$ is the same as the power of $Q^{\mathrm{perm}}_{T-1}$, even though the Besag-Clifford approach may stop before the maximum number of permutations $T$ are sampled [33]. Our sequential strategies are applied with the same parameters as described in Section 7.1.
Figure 2: Power and average number of permutations until the decision was obtained for $\alpha=0.05$ and different strengths of the alternative $\mu$ — following the experimental protocol in Subsection 7.2. Upper left plot: Empirical power; Upper right plot: Average number of permutations $\bar{\tau}$; Lower left plot: Average number of permutations until it was stopped for rejection $\bar{\tau}_{1}$; Lower right plot: Average number of permutations until it was stopped for futility $\bar{\tau}_{0}$. Relationship between the plots is given by $\bar{\tau}=(\bar{\tau_{1}}m_{1}+\bar{\tau}_{0}m_{0}+T(m-m_{1}-m_{0}))/m$, where $m_{1}$ and $m_{0}$ are the number of simulation runs where it was stopped for rejection and futility, respectively. The power of binomial strategy and binomial mixture strategy is similar as with Besag-Clifford and number of permutations is reduced substantially, particularly when the alternative is strong.
@@ 7.3 Improving power by stochastic rounding
Figure 4 shows that stochastic rounding leads to a significant improvement in power, particularly for the binomial mixture strategy in case of $\alpha=0.01$. This also implies that the binomial mixture strategy was often stopped while having a promising wealth. Therefore, drawing $1000$ permutations might be not enough when the $\alpha$ is low and the decision is difficult ($Q_{\mathrm{lim}}$ close to $\alpha$). Hence, the binomial mixture strategy is likely to achieve further rejections if we would continue sampling. In Figure 5 we compare the power of these randomized strategies with the Besag-Clifford method. There is hardly any power difference visible. In particular, the Besag-Clifford method and randomized binomial mixture strategy seem to overlap completely. Note that our methods obtained this power while reducing the number of permutations considerably (Figure 2 and Figure 3) and offering the option to stop at any point in time.
In Appendix F, we provide additional simulations results comparing the randomized binomial mixture strategy with the Besag-Clifford method in terms of power and number of permutations for different parameter configurations. All results show the same behavior, the two methods provide a similar power, while the binomial mixture strategy reduces the number of permutations substantially.
@@ 7.4 Real data: testing Fisher’s sharp null
We consider a treatment vs. control trial with binary outcomes that was already used by Ding [8] and Rosenbaum [28]. There were 18 successes among the 32 treated observations and 5 successes among the 21 control observations. We are interested in testing Fisher’s sharp null hypothesis of no individual treatment effect $H_{0}^{\mathrm{IT}}$ at level $\alpha=0.05$. Under the Stable Unit Treatment Value Assumption (SUTVA; [29]), which states that there is only one version of the treatment and no interference between the individuals, this null hypothesis can be tested using the permutation test described in Example 1.
@@
Fisher’s exact test offers a p-value of $0.026$ and thus the null hypothesis could be rejected by this test. We applied our binomial strategy (Algorithm 2) and binomial mixture strategy with density $u_{c}$, $c=0.95\alpha$, $1000$ times with a maximum number of $5000$ permutations. In this case, we omitted the stop for futility to not affect the power. The null hypothesis was rejected by both strategies in all $1000$ runs. The binomial strategy needed a mean number of $85$ permutations and a median number of $53$ permutations, while the binomial mixture strategy required a mean number of $147$ permutations and median number of $97$ permutations to obtain the decision. As a comparison, we also applied the Besag-Clifford method with $h=10$ and $T=h/\alpha$, leading to a mean (and median) number of $200$ permutations. Note that in this case the mean (and median) number of permutations is greater than with our sequential approaches. Still, the Besag-Clifford method failed to reject in $3$ of the $1000$ cases. If $h$ is reduced to $8$ (leading to a mean and median number of $160$ permutations), it even failed to reject in $16$ cases. The reason for this gain in efficiency by our methods is that they adapt the number of permutations to the difficulty of the decision, as also described in the previous subsection. Also note that the binomial mixture strategy even guarantees to reject at some finite time in this case, as the true p-value is below $c=0.0475$.
We illustrated the wealth processes of our sequential strategies for different runs and the first $500$ steps in Figure 6. Since the binomial mixture strategy converges to $1/c$, we expect it to bet more safely when it approaches $1/\alpha$. For this reason, the binomial strategy allows to reject the hypothesis earlier in most cases. However, when there are a lot of losses, e.g. as in the beginning of the upper right plot, the binomial mixture strategy leads to a larger wealth and thus seeming more promising when the evidence is not very strong. Also note that when the number of wins is much larger than the number of losses, as in the lower plots, the binomial strategy begins to lose wealth when further wins occur — after the hypothesis is already rejected. Due to (10), this is the case if $(L_{t-1}+1)/(t+1)<p$, where $p=1/\lceil\sqrt{2\pi e^{1/6}}/\alpha\rceil$.
@@ 7.5 Real data: testing conditional independence under model-X
Here, we consider testing the independence between two variables $W_{0}$ and $X$ conditional on some covariates $Z$, as described in Example 2. We follow the analysis by Grünwald et al. [12] and Berrett et al. [2] and apply our methods on the Capital bikeshare dataset, which is available at https://ride.capitalbikeshare.com/system-data. In this example, $W_{0}$ denotes the logarithm of the ride duration, $X$ the binary variable membership (casual users vs. long-term membership) and $Z$ is the three dimensional vector of starting point, destination and starting time. For the generation of the test statistics $Y_{0},Y_{1},\ldots$, we used the code of Grünwald et al. [12], which is available in the supplementary material of their paper. The distribution $W_{0}|Z$ is modeled as Gaussian $N(\mu_{Z},\sigma_{Z}^{2})$, where $\mu_{Z}$ and $\sigma_{Z}^{2}$ are estimated by a kernel regression of $W_{0}$ with the time of day as covariate, separately for each combination of starting point and destination. The test statistics are then given by $Y_{t}=|\mathrm{cor}(X,W_{t}-\mathbb{E}[W_{0}|Z])|$.
We applied our binomial strategy and binomial mixture strategy $1000$ times with the same parameters as in Section 7.4 (in particular, the type I error level was set to $0.05$). In the original code of Grünwald et al. [12], the training dataset consists of $158,741$ observations and the test data of $7,173$ observations. In this case, all $Y_{i}$, $i\geq 1$, were smaller than $Y_{0}$ such that the binomial strategy needed $44$ permutations and the binomial mixture strategy $61$ permutations to reject the null hypothesis. We also applied our methods when only the first $2000$ observations of the test data were included, thus increasing the probability of $Y_{i}$ being greater or equal than $Y_{0}$. This led to a mean permutation p-value of $0.008$ and the null hypothesis was still rejected by both sequential strategies in all $1000$ runs. The binomial strategy needed a mean number of $49$ permutations and a median number of $44$ permutations, while the binomial mixture strategy required a mean number of $83$ permutations and median number of $61$ permutations to obtain the decision. Since the evidence against $H_{0}$ is very large in this data, the classical permutation p-value and the Besag-Clifford method most likely also reject with a small number of maximum permutations $T$. However, note that we usually do not know how large the evidence in the data really is and therefore it would be a risky strategy to prespecify a small $T$. In contrast, our methods provide strong guarantees regarding the resampling risk while rejecting fast due to large evidence against the null.
@@ 8 Discussion
We have introduced a new framework for the construction of anytime-valid Monte-Carlo tests based on a simple testing by betting approach. Our general method encompasses all permutation tests based on the number of losses and particularly provides anytime-valid generalizations of the classical permutation p-value and the Besag-Clifford method. We derived the (oracle) log-optimal betting strategy under the considered alternative, given the true distribution of the limiting permutation p-value. We showed that for any given working prior on the limiting permutation p-value $Q_{\mathrm{lim}}$, the mimicked log-optimal strategy can be written as a mixture of simple binomial strategies. Based on this, we proposed a strategy with zero resampling risk after a finite number of permutations for all $q_{\mathrm{lim}}\in[0,c)$, where $c<\alpha$ can be chosen arbitrarily close to $\alpha$. Furthermore, we performed a simulation study and a real data analysis which illustrates that our anytime-valid p-values have a similar power as the classical permutation p-value and the Besag-Clifford method while the number of permutations until the decision is obtained can be substantially reduced.
@@
Binomial strategy   Reduces permutations substantially   Similar power as classical permutation        p-value in all simulation scenarios   Always rejects earlier than classical        permutation p-value, if        $Q^{\mathrm{perm}}<1/\lceil\sqrt{2\pi e^{1/6}}/\alpha\rceil\approx\alpha/e$        (Proposition 4.3)   Parameter-free (see Algorithm 2)   Cannot ensure arbitrarily        small resampling risk
@@ Acknowledgments
AR thanks Leila Wehbe for posing the question of whether permutation tests can be stopped early under null and alternative, and for stimulating practical discussions. LF thanks Timothy Barry for practical feedback on the first version of the paper. LF and AR thank Peter Grünwald for helpful comments. LF and AR thank two anonymous reviewers for useful suggestions that significantly improved the paper. LF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 281474342/GRK2224/2. AR was funded by NSF grant DMS-2310718.
Appendix A More details on related work
@@
For the particular problem of testing our null hypothesis that the observations are exchangeable, there exists no nontrivial test martingale for the canonical data filtration (generated by the observations) [25]. There are two different approaches to circumvent this issue. The first idea is to replace the data filtration by a coarser one. This means the information to be released at each step is restricted. This was used by Vovk and colleagues in a conformal prediction approach [38, 39, 36], and this is also the approach taken in the current paper (modulo some differences around randomization of ranks). The second approach is to construct an e-process which is not a test martingale, based on universal inference [41, 25]. The idea is to partition the composite null hypothesis, define a test martingale for each subset and take the infimum over these test martingales to yield an e-process. In terms of testing by betting, it could be interpreted as playing multiple games, where the gambler’s wealth is the minimum wealth across all these games. In this case, the wealth process is adapted to the canonical data filtration, has expectation of at most one under the null at any stopping time in the original filtration of the data.
@@
Very recently, other sequential tests of exchangeability were proposed, based on the former approach. Saha and Ramdas [30] introduced a test martingale for Markovian alternatives that is based on “pairwise betting”, which means that two observations are observed and bet on simultaneously. Independently, Koning [17] introduced a sequential test for general group invariance (including exchangeability). In particular, he derives likelihood ratios for group invariance and defines a test martingale as running product of these. Similarly, Lardy and Pérez-Ortiz [20] construct general tests of group invariance based on conformal prediction. Our test martingale is also based on coarsening the data filtration by not revealing the $Y_{t}$ itself, but the loss indicator $I_{t}$, and is thus somewhat related to this line of work. The main difference of our approach to the previously mentioned approaches is the alternative we test against. Vovk [36] focuses on change-point alternatives, Ramdas et al. [25] and Saha and Ramdas [30] on Markovian alternatives, while Koning [17] considers general group invariance versus normal distributions with a location shift. In contrast, our paper focuses on the usual setting of permutation tests (where permutations rather than data are sampled sequentially), where under the alternative, only the original test statistic is special while all the permuted statistics are exchangeable by design. It should be noted that the above works give the mentioned applications as examples of their more general frameworks. The general frameworks may include (a modified version of) our approach as a special case, as described in Remark 3 for Vovk’s conformal prediction approach. However, none of these works make the connection to sequential permutation tests (where not the data but the permutations are drawn sequentially). We demonstrated in this paper that this special case is of great practical interest, simplified these general frameworks for our particular alternative to make the method accessible to a broader audience and constructed powerful betting strategies yielding test martingales with easy calculable closed forms.
Appendix B Calibrating $Q^{\mathrm{perm}}_{T}$ into an e-value